OpenStudy (anonymous):

Solve the system 2x - 3y = 7 and 8x - 12y = 5

OpenStudy (anonymous):

Multiply both sides of the first equation by 4 and you get, 8x-12y=28 The second equation is 8x-12y=5, So I would say this system is unsolvable.

hero (hero):

Not solveable: -15 + 24x = -84 + 24x 99 = 0x ---which isn't true

OpenStudy (anonymous):

there is no solution